alexmarrow:

an object of a mass 1.00kg is observed to have an acceleration with a magnitude of 10.0m/s^2 in a direction 60.0 degree east of north.the force f2 acting on the object has a magnitude 5.00N and its directed north.determine the magnitude and direction of the one other horizontal force.

supie:

Is there any answer choices?

10kbands:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Is there any answer choices? \(\color{#0cbb34}{\text{End of Quote}}\) yeah i'm looking for answer choices

supie:

Oh ok.

alexmarrow:

there's no answer choices sorry..please help me

SemiDefinite:

\(\vec a = 10 \left[ \sin(60), \cos(60) \right]^T \) \(\vec F_2 = \left[ 0,5 \right]^T \) let: \(\vec F_1 = \left[ x,y \right]^T \) \( m \vec a = \vec F_1 + \vec F_2 = [x, y + 5]^T = 10 \left[ \sin(60), \cos(60) \right]^T \) \(\therefore\) \(x = 5 \sqrt{3}\) \(y = 0\)